Three charges are fixed to an xy-coordinate system. A charge of 18µC is on y = 3.0m. A charge of −12µC is at the origin. Lastly, a charge of 45µC is on the x-axis at x = 3.0m. Determine the magnitude and direction of the net electrostatic force on the charge at x = 3.0m


1
Expert's answer
2021-06-30T10:19:37-0400

Let the charge of 18 µC be "q_1", the charge of −12 µC be "q_2" and the charge of 45 µC be "q_3". Let's first find the magnitude of the force with which the charge "q_1" acts on charge "q_3":


"F_{13}=\dfrac{kq_1q_3}{r_{13}^2}=\dfrac{kq_1q_3}{(\sqrt{r_{12}^2+r_{23}^2})^2},""F_{13}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot18\cdot10^{-6}\ C\cdot45\cdot10^{-6}\ C}{(\sqrt{(3\ m)^2+(3\ m)^2})^2}=0.405\ N."

Similarly, let's find the magnitude of the force with which the charge "q_2" acts on charge "q_3":


"F_{23}=\dfrac{kq_2q_3}{r_{23}^2},""F_{23}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot12\cdot10^{-6}\ C\cdot45\cdot10^{-6}\ C}{(3\ m)^2}=0.54\ N."

Let's find the projections of forces "F_{13}" and "F_{23}" on axis "x" and "y":


"F_{13,x}=F_{13}cos45^{\circ}=0.405\ N\cdot cos45^{\circ}= 0.286\ N,""F_{13,y}=F_{13}sin45^{\circ}=0.405\ N\cdot sin45^{\circ}= -0.286\ N,""F_{23,x}=-0.54\ N,""F_{23,y}=0."

Then, we get:


"F_x=F_{13,x}+F_{23,x}=0.286\ N+(-0.54\ N)=-0.254\ N,""F_y=F_{13,y}=-0.286\ N."

We can find the net electrostatic force on the charge "q_3" from the Pythagorean theorem:


"F=\sqrt{F_x^2+F_y^2},""F=\sqrt{(-0.254\ N)^2+(-0.286\ N)^2}=0.382\ N."

We can find the direction of the net electrostatic force on the charge "q_3" from the geometry:


"tan\theta=\dfrac{F_y}{F_x},""\theta=tan^{-1}(\dfrac{F_y}{F_x}),""\theta=tan^{-1}(\dfrac{-0.286\ N}{-0.254\ N})=48.4^{\circ}."

The net electrostatic force on the charge "q_3" directed at "48.4^{\circ}" below the "x"-axis.